OpenStudy (anonymous):

how do you solve v^2+6v-59=0 by completing the square?

OpenStudy (anonymous):

\[v^2+6v=59\] \[(v+3)^2=59+9=68\] \[v+3=\pm\sqrt{68}\] \[v=-3\pm\sqrt{68}\]

OpenStudy (anonymous):

or if you prefer, \(v=-3\pm2\sqrt{17}\)